# Tensorization of neural networks for improved privacy and interpretability

### v2

@@ 3.2. Neural networks
In both cases, pivots are selected from the corresponding training sets, with $N$ ranging from 30 to 100. The test samples used to measure accuracy are taken from the test sets, fixing $M=1000$, as in previous experiments. Additionally, all the TTs are set to have a fixed rank $r=10$, with $c=1-10^{-5}$.
[…]
Figure 3.2. Performance of TT-RSS applied to different NN models when varying the number of pivots $N$. From left to right, columns show: relative error on the $N$ pivots $\mathbf{x}$ used in TT-RSS, $R(\mathbf{x})$; relative error on a set of $M$ test samples $\mathbf{s}$ from the test sets, $R(\mathbf{s})$; percentage of classifications of the TT models that differ from those of the original NN models; and time in seconds taken to perform the decompositions. For each value of $N$, the decomposition is performed 10 times, displaying in the figures the mean values with error bars at ±$\pm 0.5\sigma$, where $\sigma$ denotes the standard deviation. Also, three configurations are considered with different numbers of variables, $n=144$, $n=256$, and $n=400$, which are displayed using different colors.
@@ 3.2.2. MNIST
Even though this dataset can still be considered a toy dataset, it allows us to demonstrate the performance of the obtained TT models in a slightly more realistic scenario. In this case, each pixel is not binary but is considered as a continuous value between 0 and 1. Thus, a non-trivial continuous embedding is required for the input variables. Inspired by the use of TTs as approximations of polynomials, as presented in Ref. [41], we use the embedding $\phi^{i}(x)=x^{(i-1)}$, for $i\in[d]$, which we refer to as the polynomial embedding. Using these functions, the TT models represent polynomials of the variables $x_{1},\ldots,x_{n}$ up to degree $d-1$ in each variable, where the coefficients are given by the contraction of the TT. This is, the vector of coefficients is represented as a TT. In this experiment, we keep $d=2$, as in the previous experiments, which allows us to represent polynomials of degree 1 in each variable. However, in this case, the dimension of the output variable $y$ is 10, corresponding to the number of classes in the dataset. Thus, for the position in the middle of the TT where the output variable is placed, we have $d_{\lfloor n/2\rfloor}=10$. Consequently, since this variable is discrete, we use the trivial embedding for that position.
@@
We observe that the TT models achieve reasonable accuracies between 75% and 80%, taking only from 20 to 60 seconds in many cases. While this is far from what can be achieved through gradient descent optimization, the results from TT-RSS can serve as starting points for the optimization process. To achieve even better accuracies, one could also increase the allowed rank $r$ and use more pivots $N$, which would correspondingly increase the computation time.
As mentioned earlier, the low errors encountered on pivots might lead to overfitting, which could also raise concerns about generalization and privacy. However, two points must be considered: (i) when forming the tensors $p(\mathbf{x}_{1:k-1},\mathbf{y}_{k},\mathbf{x}_{k+1:n})$, $dN^{2}$ points are used to tensorize $p$ (in contrast with the $N$ pivots), which helps improve generalization, and (ii) having low errors on output values does not directly imply overfitting in the usual sense. It simply indicates that the TT reproduces the output probabilities given by the NN in those cases. However, as we observe in the experiments, and as presented in the experiments of Section 4.1 (specifically, in Fig. 4.2), accuracies on the set of pivots are usually similar to the accuracies on test samples.
@@ 3.3. Dependence on hyperparameters
In the previous experiments, it was observed that when repeating the same tensorizations multiple times, the results could vary significantly in some situations. This is due to the inherent randomness introduced in TT-RSS, including the random selection of the pivots and the random matrices $U_{k}$. In this section, we tensorize a single NN model multiple times, varying the hyperparameters of the algorithm—$d$, $r$, and $N$—to examine how they affect accuracy and the variability of results.
@@
(3.7) $\text{acc}(p,\mathbf{s},y)=\frac{1}{M}\sum_{i=1}^{M}\mathcal{X}_{\{\arg\max_{y}p(y|\mathbf{s}^{i})\}}(y^{i}),$
where $\mathcal{X}$ denotes the indicator function, $\mathbf{s}$ is the set of test samples, and $y$ is the vector of labels for the samples. Results are presented in Fig. 3.3.
Figure 3.3. Distribution of the accuracies of 50 tensorized models for each configuration, represented using box plots. Each configuration is made by varying one of the hyperparameters $d$, $r$, and $N$ from the baseline point $d=2$, $r=5$, $N=50$. The orange lines connect the medians across all the different values for a hyperparameter. The horizontal dashed black lines represent the original accuracy of the NN model being tensorized.
@@ 4.1. Privacy
To address this limitation, we present a much more realistic scenario in this section, where the objective is to classify voices by gender. In this case, each voice sample is represented as a vector in $[0,1]^{500}$, and the irrelevant feature targeted for the attack is one that would not typically be removed during standard pre-processing and does not correspond to a single input variable: the accent. To demonstrate the privacy leakage, we train NN models via gradient descent using training datasets where the accents are predominantly of one type or another. As a solution, we show that instead of training TT models from scratch, TT-RSS can be used to tensorize the original models, reconstructing models that are private while maintaining nearly the same accuracies. In cases where the accuracies drop, it is possible to re-train the TT models for a few additional epochs. In either case, this approach is more efficient and straightforward than training new TT models from the ground.
@@ 4.1.1. Datasets
To enable statistical analysis, we partition the dataset into several subsets, training multiple NNs on the different subsets. The partitions are also stratified based on the proportion of the accent variable. First, we extract a subset from the predefined train partition of CommonVoice and balance the gender variable across the accents. Denoting $q$ as the percentage of English speakers, we form, for each $q\in\{5,10,20,30,40,50,60,70,80,90,95\}$, 10 datasets. For each value of $q$, we create an imbalanced dataset according to the specified accent proportion, and sample 80% of the client_id identifiers to form the 10 datasets, splitting each of them into training and validation subsets. Consequently, the 10 datasets created for each $q$ value are distinct but not disjoint. From the remaining 20% of IDs in each case, we select around 200 samples to serve as the pool from which the $N$ pivots for tensorizing the NN models are chosen. Thus, the pivots used in TT-RSS exhibit the same imbalances as those present in the corresponding training datasets. These subsets of 200 samples will be used also to re-train the TT models.
For pre-processing the audio samples, we first resample all audios to a frequency of 1000 Hz and crop each to retain only the middle 1000 values, corresponding to one second of audio per person. Next, we apply the Fast Fourier Transform (FFT) and retain only half of the values, since the FFT of real-valued functions is Hermitian-symmetric. This results in feature vectors with 500 variables. Finally, we renormalize the values to restrict them to the interval $[0,1]$.
@@ 4.1.2. Models
Using these hyperparameters, we train 25 models for 20 epochs on each of the 10 datasets corresponding to different imbalance levels $q$. The resulting models achieve an average accuracy of approximately 85%.
@@ 4.1.3. Tensorization
To obfuscate the TT cores, we introduce random orthogonal (in general, unitary) matrices $W_{k}$ and their transposes (equivalently, adjoints), $W_{k}^{\intercal}$, between contractions. This results in new cores defined as
(4.2) $\widetilde{G}_{k}(\alpha_{k-1},x_{k},\alpha_{k}):=W_{k-1}^{\intercal}(\alpha_{k-1},\theta_{k-1})G_{k}(\theta_{k-1},x_{k},\theta_{k})W_{k}(\theta_{k},\alpha_{k}).$
@@
Figure 4.2. Difference between the accuracy measured only on the pivots used in TT-RSS and the accuracy on test samples. On the left, accuracies are measured for TT models directly output from the TT-RSS decomposition. On the right, accuracies are for TT models after re-training them for 10 epochs on a small dataset containing the pivots. These differences are measured for all percentages $q$ of the presence of English speakers in the datasets. For each $q$, mean differences in accuracy across all models trained for that $q$ are shown, with error bars at ±$\pm 0.5\sigma$, where $\sigma$ denotes the standard deviation.
@@ 4.1.4. Attacks
As detailed earlier, we trained 250 NNs for each percentage $q$. To train the attack model, we merge all models corresponding to $q$ and $100-q$, for $q\in\{5,10,20,30,40\}$, and create a dataset by storing some information from each model, $x$, along with its corresponding label, $y$. We define the label for each model as $y=0$ if $q<50$, and $y=1$ otherwise. For the information representing the model, $x$, we consider two scenarios: (i) black-box access, where we can only observe classifications made by the model, and (ii) white-box access, where we have full access to the model’s parameters.
@@
Figure 4.3. Accuracy of logistic regression models trained to classify whether a model was trained on a dataset with a majority of English speakers or a majority of Canadian speakers. For each percentage of imbalance, a dataset is formed from 500 models trained on 20 different datasets (10 datasets per percentage $q$, with 25 models per dataset). For each configuration, attacks are evaluated via $K$-fold validation with $K=10$, and mean accuracies are displayed with error bars at ±$\pm 0.5\sigma$, where $\sigma$ denotes the standard deviation.
@@ 4.2. Interpretability
In recent years, several approaches have emerged to approximate quantum states via neural network representations, enabling the efficient representation of complex systems [13, 14, 15, 16, 17]. However, these representations often lack the capacity to analyze properties of the states directly, requiring additional models to estimate each property separately [18, 19, 20]. Conversely, the methodology we propose leverages pre-trained NNQS to derive a TN representation, enabling the direct study of such properties.
@@ 4.2.1. AKLT Model
We present this methodology in a simplified scenario that may serve as a proof of principle. In particular, we will estimate the SPT phase of the AKLT state [67]. Although this state admits a TT form, this is not necessary for the method to succeed, since we will only allow black-box access to this function. Indeed, we will exploit the TT structure of the state only when obtaining the pivots and only for computational efficiency. Thus, this setup is analogous to having access to other less interpretable models, such as a NN, which we can query with spin configurations to obtain the associated amplitudes. Using this function, we will reconstruct a new TT representation of the state via TT-RSS and use this representation to compute the order parameter defined in Ref. [68].
@@
(4.3) $H=\sum_{i=1}^{n}P^{(2)}_{i,i+1},$
where $P^{(2)}_{i,i+1}$ is the projector operator that projects the total spin of the neighboring sites $i$ and $i+1$, both of them spin-1 particles, onto the spin-2 subspace. Consequently, its ground state, referred to as the AKLT state, is formed by spin-1/2 singlets between nearest neighbor sites, with projector operators at each site projecting the two spin-1/2’s into the spin-1 subspace.
This state can be represented in TT form, or as commonly denoted in the physics literature, as a MPS. For the representation of the model and the subsequent description of the order parameter computation, we use Dirac’s bra-ket notation. Additionally, instead of representing the TT cores as tensors $G_{k}(\alpha_{k-1},x_{k},\alpha_{k})$, we will simply denote them as matrices indexed by the input label, $G_{k}^{x_{k}}$. Thus, a translationally invariant TT representation of the AKLT state is defined as:
(4.4) |$\ket{\Psi}=\sum_{x_{1},\ldots,x_{n}}\mathrm{Tr}[G^{x_{1}}\cdots G^{x_{n}}]\ket{x_{1}\cdots x_{n}},$

### v3

@@ 3.2. Neural networks
In both cases, pivots are selected from the corresponding training sets, with $N$ ranging from 30 to 100. The test samples used to measure accuracy are taken from the test sets, fixing M=1  [^0] M=1\,000, as in previous experiments. Additionally, all the TTs are set to have a fixed rank $r=10$, with $c=1-10^{-5}$.
Finally, apart from measuring the relative errors, we evaluate how different the reconstructed TT models are from the original NNs by measuring, over the $M$ test samples, the percentage of classifications given by the TTs that differ from those given by the NNs. This is a simple metric that indicates, in a more faithful manner than the relative error, how similarly both models perform in practice. Although this does not directly measure the accuracy of the TT models, in these particular examples, where the accuracy of the NN models is close to 100% (100% test accuracy on Bars and Stripes, and  98% on MNIST), the accuracy of the TTs can be approximated as 100(percentage of different classifications)100-$\text${(percentage of different classifications)}. The results are depicted in Fig. 3.2.
@@ 3.2.1. Bars and Stripes
Figure 3.2. Performance of TT-RSS applied to different NN models when varying the number of pivots NN (x-axis). From left to right, columns show: relative error on the NN pivots $\mathbf{x}$ used in TT-RSS, $R(\mathbf{x})$; relative error on a set of $M$ test samples $\mathbf{s}$ from the test sets, $R(\mathbf{s})$; percentage of classifications of the TT models that differ from those of the original NN models; and time in seconds taken to perform the decompositions. For each value of $N$, the decomposition is performed 10 times, displaying in the figures the mean values with error bars at $\pm 0.5\sigma$, where $\sigma$ denotes the standard deviation. Also, three configurations are considered with different numbers of variables, $n=144$, $n=256$, and $n=400$, which are displayed using different colors. All computations are performed in double precision.
@@ 3.2.2. MNIST
Even though this dataset can still be considered a toy dataset, it allows us to demonstrate the performance of the obtained TT models in a slightly more realistic scenario. In this case, each pixel is not binary but is considered as a continuous value between 0 and 1. Thus, a non-trivial continuous embedding is required for the input variables. Inspired by the use of TTs as approximations of polynomials, as presented in Ref. [41], we use the embedding $\phi^{i}(x)=x^{(i-1)}$, for $i\in[d]$, which we refer to as the polynomial embedding. Using these functions, the TT models represent polynomials of the variables $x_{1},\ldots,x_{n}$ up to degree d1d-1 in each variable, where the coefficients are given by the contraction of the TT. This is, the vector of coefficients is represented as a TT. In this experiment, we keep $d=2$, as in the previous experiments, which allows us to represent polynomials of degree 1 in each variable. However, in this case, the dimension of the output variable $y$ is 10, corresponding to the number of classes in the dataset. Thus, for the position in the middle of the TT where the output variable is placed, we have $d_{\lfloor n/2\rfloor}=10$. Consequently, since this variable is discrete, we use the trivial embedding for that position.
@@
We observe that the TT models achieve reasonable accuracies between 75% and 80%, taking only from 20 to 60 seconds in many cases. While this is far from what can be achieved through gradient descent optimization, the results from TT-RSS can serve as starting points for the optimization process. To achieve even better accuracies, one could also increase the allowed rank $r$ and use more pivots NN, which would correspondingly increase the computation time.
As mentioned earlier, the low errors encountered on pivots might lead to overfitting, which could also raise concerns about generalization and privacy. However, two points must be considered: (i) when forming the tensors $p(\mathbf{x}_{1:k-1},\mathbf{y}_{k},\mathbf{x}_{k+1:n})$, $dN^{2}$ points are used to tensorize $p$ (in contrast with the NN pivots), which helps improve generalization, and (ii) having low errors on output values does not directly imply overfitting in the usual sense. It simply indicates that the TT reproduces the output probabilities given by the NN in those cases. However, as we observe in the experiments, and as presented in the experiments of Section 4.1 (specifically, in Fig. 4.2), accuracies on the set of pivots are usually similar to the accuracies on test samples.
@@
3.2.3. Comparison with TT-CI
As in the case of the non-densities, we also compare the performance of TT-RSS and TT-CI for the tensorization of NN models. The results, including relative test errors, percentage of differing classifications, and time costs, are shown in Table 3.2. As before, times in parentheses for TT-CI denote the total time taken by the full algorithm, including preliminary steps. For TT-RSS, the reported time already includes the entire process, consisting of a single left-to-right sweep. In TT-CI, we also fix the ranks in advance to $r=10$, and set the number of iterations—one left-to-right and one right-to-left sweep—to 1 for the Bars and Stripes dataset, and 5 for MNIST.
For the Bars and Stripes dataset, we observe very similar test errors, with the most notable difference being in time costs. In this case, both methods yield comparable errors and classification differences, while TT-RSS is faster, especially when considering the total time of TT-CI. Notably, for $n=400$, we observe a larger difference in accuracy, with TT-CI performing better. This may indicate that a larger pivot set is needed to improve the accuracy of TT-RSS in this higher-dimensional model.
The most relevant result for our purposes—and the one that most clearly supports our approach—is the significant loss in accuracy observed for MNIST when using TT-CI. The only configuration we consider for MNIST is $n=144$, the simplest case. Moreover, we allow TT-CI a larger number of iterations to enable a more extensive search over the model’s domain in order to find suitable pivots. Nevertheless, the results show a test error close to 1, suggesting that the TT approximation produced by TT-CI may output values close to 0 across all test points. This is reflected in a classification difference of $84.1\%$, indicating a completely erroneous model. In contrast, TT-RSS with $N=100$ yields larger test errors, but a classification difference of only $20.7\%$, implying that the model has around $80\%$ accuracy. These results highlight how, for more intricate distributions—typically flat with narrow peaks—the maxvol routine fails to identify a suitable set of pivots that adequately captures the model’s behavior. In such cases, incorporating additional knowledge about the model’s domain becomes crucial, as already hinted in Ref. [77].
However, although TT-RSS performs significantly better, its results are still far from optimal. As discussed earlier, training the TT model via gradient descent can achieve accuracies close to the original $98\%$. Yet this leads to a different model, providing a distinct approximation of the underlying distribution. Our goal through tensorization is to preserve the form of the original NN approximation within a TT representation. This may lead to larger errors, which in turn affect accuracy. This limitation is especially pronounced for MNIST, whose 2D structure, and the spatial patterns likely learned by the original NN, do not align well with the TT format. In contrast, other models, such as the audio classification model discussed in Section 4.1, do not show such drops in accuracy (see Fig. 4.1).
Finally, we note that using double precision in these examples does not affect accuracy. Only in cases where the function has an exact TT representation, such as in the random TT examples, do errors reach machine precision, and double precision can enhance results. For all other functions and NN models, where errors remain far from machine precision, there is no observable difference between single and double precision, aside from a noticeable increase in time cost when using double precision. For this reason, we use double precision only in Section 4.2, and single precision elsewhere.
Table 3.2. Performance comparison of TT-RSS and TT-CI applied to different NN models with varying numbers of variables $n$ and pivots NN. The table reports the relative error on a set of $M$ test samples $\mathbf{s}$ from the test sets, $R(\mathbf{s})$; the percentage of classifications of the TT models that differ from those of the original NN models; and the time (in seconds) required to perform the decompositions. For TT-CI, the reported time corresponds to the sweeps, while the value in parentheses includes the total time of the full algorithm, including preliminary steps for decomposition setup. Each decomposition is repeated 10 times, and the median values are shown. All computations are performed in double precision.
  $f$ $n$ Tensorization algorithm Error on test % of diff. classifications Time in secs.
  $T$-RSS ($N=50$) $5.97\times 10^{-1}$ $0.9\%$ $7.03\times 10^{0}$
144 $T$-RSS ($N=100$) $3.95\times 10^{-1}$ $0.5\%$ $2.29\times 10^{1}$
$T$-CI $4.07\times 10^{-1}$ $2.1\%$ $4.23\times 10^{1}$ ($5.65\times 10^{2}$)
$T$-RSS ($N=50$) $4.12\times 10^{-1}$ $1.2\%$ $1.49\times 10^{1}$
Bars and Stripes 256 $T$-RSS ($N=100$) $4.52\times 10^{-1}$ $5\%$ $4.96\times 10^{1}$
$T$-CI $4.39\times 10^{-1}$ $1.5\%$ $8.62\times 10^{1}$ ($1.70\times 10^{3}$)
$T$-RSS ($N=50$) $5.08\times 10^{-1}$ $13.7\%$ $2.54\times 10^{1}$
400 $T$-RSS ($N=100$) $5.09\times 10^{-1}$ $15.1\%$ $8.69\times 10^{1}$
$T$-CI $5.07\times 10^{-1}$ $3.9\%$ $1.55\times 10^{2}$ ($4.35\times 10^{3}$)
$T$-RSS ($N=50$) $1.17\times 10^{1}$ $31.2\%$ $8.26\times 10^{0}$
MNIST 144 $T$-RSS ($N=100$) $5.51\times 10^{0}$ $20.7\%$ $2.44\times 10^{1}$
$T$-CI $1.20\times 10^{0}$ $84.1\%$ $2.12\times 10^{2}$ ($7.37\times 10^{2}$)
@@ 3.3. Dependence on hyperparameters
In the previous experiments, it was observed that when repeating the same tensorizations multiple times, the results could vary significantly in some situations. This is due to the inherent randomness introduced in TT-RSS, including the random selection of the pivots and the random matrices $U_{k}$. In this section, we tensorize a single NN model multiple times, varying the hyperparameters of the algorithm—$d$, $r$, and NN—to examine how they affect accuracy and the variability of results.
@@
(3.7) $\text{acc}(p,\mathbf{s},\mathbf{y})=\frac{1}{M}\sum_{i=1}^{M}\chi_{\{\arg\max_{y}p(y|\mathbf{s}^{i})\}}(\mathbf{y}^{i}),$
where $\chi$ denotes the indicator function, $\mathbf{s}$ is the set of test samples, and $\mathbf{y}$ is the vector of labels for the samples. Results are presented in Fig. 3.3.
Figure $3.$ Distribution of the accuracies of 50 tensorized models for each configuration, represented using box plots. Each configuration is made by varying one of the hyperparameters $d$, $r$, and NN from the baseline point $d=2$, $r=5$, $N=50$. The orange lines connect the medians across all the different values for a hyperparameter. The horizontal dashed black lines represent the original accuracy of the NN model being tensorized.
@@ 4.1. Privacy
To address this limitation, we present a much more realistic scenario in this section, where the objective is to classify voices by gender. In this case, each voice sample is represented as a vector in $0,1]^{500}$, and the irrelevant feature targeted for the attack is one that would not typically be removed during standard pre-processing and does not correspond to a single input variable: the accent. To demonstrate the privacy leakage, we train NN models via gradient descent using training datasets where the accents are predominantly of one type or another. As a solution, we show that instead of training TT models from scratch, TT-RSS can be used to tensorize the original models, reconstructing models that are private while maintaining nearly the same accuracies. In cases where the accuracies drop, it is possible to re-train the TT models for a few additional epochs. In either case, this approach is more efficient and straightforward than training new TT models from the ground.
@@ 4.1.1. Datasets
To enable statistical analysis, we partition the dataset into several subsets, training multiple NNs on the different subsets. The partitions are also stratified based on the proportion of the accent variable. First, we extract a subset from the predefined train partition of CommonVoice and balance the gender variable across the accents. Denoting $q$ as the percentage of English speakers, we form, for each $q\in\{1,5,10,20,30,40,50,60,70,80,90,95,99\}$, 10 datasets. For each value of $q$, we create an imbalanced dataset according to the specified accent proportion, and sample 80% of the client_id identifiers to form the 10 datasets, splitting each of them into training and validation subsets. Consequently, the 10 datasets created for each $q$ value are distinct but not disjoint. From the remaining 20% of IDs in each case, we select around 200 samples to serve as the pool from which the NN pivots for tensorizing the NN models are chosen. Thus, the pivots used in TT-RSS exhibit the same imbalances as those present in the corresponding training datasets. These subsets of 200 samples will be used also to re-train the TT models.
For pre-processing the audio samples, we first resample all audios to a frequency of 1 0001\,000 Hz and crop each to retain only the middle 1 0001\,000 values, corresponding to one second of audio per person. Next, we apply the Fast Fourier Transform (FFT) and retain only half of the values, since the FFT of real-valued functions is Hermitian-symmetric. This results in feature vectors with 500 variables. Finally, we renormalize the values to restrict them to the interval $[0,1]$.
@@ 4.1.2. Models
Using these hyperparameters, we train 25 models for 20 epochs on each of the 10 datasets corresponding to different imbalance levels $q$. The resulting models achieve an average accuracy of approximately 82%.
@@ 4.1.3. Tensorization
To obfuscate the TT cores, we introduce random orthogonal (in general, unitary) matrices $W_{k}$ and their transposes (equivalently, adjoints), $W_{k}^{\top}$, between contractions. This results in new cores defined as
(4.2) $\widetilde{G}_{k}(\alpha_{k-1},x_{k},\alpha_{k}):=W_{k-1}^{\top}(\alpha_{k-1},\theta_{k-1})G_{k}(\theta_{k-1},x_{k},\theta_{k})W_{k}(\theta_{k},\alpha_{k}).$
@@
Figure 4.2. Difference between the accuracy measured only on the pivots used in TT-RSS and the accuracy on test samples. On the left, accuracies are measured for TT models directly output from the TT-RSS decomposition. On the right, accuracies are for TT models after re-training them for 10 epochs on a small dataset containing the pivots. These differences are measured for all percentages $q$ of the presence of English speakers in the datasets. For each $q$, mean differences in accuracy across all models trained for that $q$ are shown, with error bars at $\pm 0.5\sigma$, where $\sigma$ denotes the standard deviation.
@@ 4.1.4. Attacks
As detailed earlier, we trained 250 NNs for each percentage $q$. To train the attack model, we merge all models corresponding to $q$ and 100q100-q, for $q\in\{1,5,10,20,30,40\}$, and create a dataset by storing some information from each model, $x$, along with its corresponding label, $y$. We define the label for each model as $y=0$ if $q<50$, and $y=1$ otherwise. For the information representing the model, $x$, we consider two scenarios: (i) black-box access, where we can only observe classifications made by the model, and (ii) white-box access, where we have full access to the model’s parameters.
@@
Figure 4.3. Accuracy of logistic regression models trained to classify whether a model was trained on a dataset with a majority of English speakers or a majority of Canadian speakers. For each percentage of imbalance, a dataset is formed from 500 models trained on 20 different datasets (10 datasets per percentage $q$, with 25 models per dataset). For each configuration, attacks are evaluated via $K$-fold validation with $K=10$, and mean accuracies are displayed with error bars at $\pm 0.5\sigma$, where $\sigma$ denotes the standard deviation.
@@ 4.2. Interpretability
In recent years, several approaches have emerged to approximate quantum states via neural network representations, enabling the efficient representation of complex systems [13, 14, 15, 16, 17]. However, these representations often lack the capacity to analyze properties of the states directly, requiring additional models to estimate each property separately [18, 19, 20]. In contrast, the methodology we propose leverages pre-trained NNQS to derive a TN representation, enabling the direct study of such properties.
@@ 4.2.1. AKLT Model
We present this methodology in a simplified scenario that may serve as a proof of principle. In particular, we will estimate the SPT phase of the AKLT state [67], which admits a TT representation that was already derived in early works [115, 116]. Although this state admits a TT form, this is not necessary for the method to succeed, since we will only allow black-box access to this function. Indeed, we will exploit the TT structure of the state only when obtaining the pivots and only for computational efficiency. Thus, this setup is analogous to having access to other less interpretable models, such as a NN, which we can query with spin configurations to obtain the associated amplitudes. Using this function, we will reconstruct a new TT representation of the state via TT-RSS and use this representation to compute the order parameter defined in Ref. [68].
@@
(4.3) $H=\sum_{k=1}^{n}P^{(2)}_{k,k+1}$,
where $P^{(2)}_{k,k+1}$ is the projector operator that projects the total spin of the neighboring sites $k$ and $k+1$, both of them spin-1 particles, onto the spin-2 subspace. Consequently, its ground state, referred to as the AKLT state, is formed by spin-1/2 singlets between nearest neighbor sites, with projector operators at each site projecting the two spin-1/2’s into the spin-1 subspace.
This state can be represented in TT/MPS form. For the representation of the model and the subsequent description of the order parameter computation, we use Dirac’s bra-ket notation. Additionally, instead of representing the TT cores as tensors $G_{k}(\alpha_{k-1},x_{k},\alpha_{k})$, we will simply denote them as matrices indexed by the input label, $G_{k}^{x_{k}}$. Thus, a translationally invariant TT representation of the AKLT state is defined as:
[…]